Retangle diagonal length is 50ft. Perimeter is 140 feet. What are the dimensions?
Answer by kapilsinghi(68)

sum of 2 adjesent sides = 140/2 = 70
let one side = x
other side = 70 - x
50^2 = (x^2 + (70-x)^2)
2500 = x^2 + 4900 +x^2 -140x
2x^2 + 4900 - 140x - 2500 = 0
2x^2 - 140x + 2400 = 0
x^2 - 70x + 1200 = 0
x^2 - 30x - 40x + 1200 = 0
x(x - 30) - 40(x + 30) = 0
(x - 40)(x - 30) = 0
x = 40 or x =30
one side = 40
other side = 30
kapilsinghi123@gmail.com